# On Closed-Form Expressions for the Fisher–Rao Distance

### v2

@@ 5.9 Power function
    which is negative, but not constant, where $\psi^{(m)}(x)\coloneqq\frac{\mathrm{d}^{m+1}}{\mathrm{d}x^{m+1}}\log\Gamma(x)$ denotes the polygamma function of order $m$. Bounds for the Fisher–Rao distance in this manifold have been studied in [15].
@@
    $\kappa$ $=\frac{\psi^{(2)}(\alpha)\psi^{(2)}(\beta)\psi^{(2)}(\alpha+\beta)}{4\left(\psi^{(1)}(\alpha)\psi^{(1)}(\beta)-\psi^{(1)}(\alpha+\beta)[\psi^{(1)}(\alpha)+\psi^{(1)}(\beta)]\right)^{2}}\left(\frac{\psi^{(1)}(\alpha)}{\psi^{(2)}(\alpha)}+\frac{\psi^{(1)}(\beta)}{\psi^{(2)}(\beta)}-\frac{\psi^{(1)}(\alpha+\beta)}{\psi^{(2)}(\alpha+\beta)}\right)$
@@
    which too is negative, but not constant. In fact, more generally, the sectional curvature of the statistical manifold of Dirichlet distributions (which are the multivariate generalisation of Beta distributions) is negative [40, Thm. 6].
@@ 5.10 Wishart
defined for $X\in P_{m}(\mathbb{R})$, and characterised by $\Sigma\in P_{m}(\mathbb{R})$, where $\Gamma_{m}(z)\coloneqq\pi^{m(m-1)/4}\prod_{j=1}^{m}\Gamma(z+(1-j)/2)$ denotes the multivariate Gamma function. Note that, for fixed $m,n$, the associated $m(m+1)/2$-dimensional statistical manifold $\mathcal{S}$ is in correspondence with the cone $P_{m}(\mathbb{R})$ of symmetric positive-definite matrices via the bijection $\iota\colon P_{m}(\mathbb{R})\to\mathcal{S}$, given by $\Sigma\mapsto\frac{{(\det X)}^{(n-m-1)/2}\exp\left(-\frac{1}{2}\operatorname{tr}(\Sigma^{-1}X)\right)}{2^{nm/2}(\det\Sigma)^{n/2}\,\Gamma_{m}(n/2)}$. Denoting $\sigma_{i,j}$ the $(i,j)$-th entry of the matrix $\Sigma$, we can write the parameter vector as $\xi=\left(\xi^{1},\dots,\xi^{m(m+1)/2}\right)=(\sigma_{1,1},\dots\sigma_{1,m},\sigma_{2,2},\dots,\sigma_{2,m},\dots,\sigma_{m,m})$. We then have $\partial_{i}\ell\coloneqq\partial_{{i}}\ell(\Sigma)=\frac{1}{2}\operatorname{tr}\left(\Sigma^{-1}X\Sigma^{-1}{(\partial_{i}\Sigma)}\right)-\frac{n}{2}\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma)}\right)$ and $\partial_{j}\partial_{i}\ell=-\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma)}\Sigma^{-1}{(\partial_{j}\Sigma)}\Sigma^{-1}X\right)+\frac{n}{2}\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma)}\Sigma^{-1}{(\partial_{j}\Sigma)}\right)$, where the derivative in $\partial_{i}\Sigma$ is done entry-wise. The elements of the Fisher metric are then
$g_{ij}(\xi)$ $=-\mathbb{E}\left[\partial_{j}\partial_{i}\ell\right]=-\mathbb{E}\left[-\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma)}\Sigma^{-1}{(\partial_{j}\Sigma)}\Sigma^{-1}X\right)-\frac{n}{2}\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma})\Sigma^{-1}{(\partial_{j}\Sigma)}\right)\right]$
@@
In view of the bijection $\iota$, the tangent space $T_{p_{\xi}}\mathcal{S}$ can be identified with $H_{m}$, the set of $m\times m$ symmetric matrices [11, Ch. 6]. Given two matrices $U,\ \widetilde{U}\in H_{m}$, parametrized as $\theta=\left(\theta^{1},\dots,\theta^{m(m+1)/2}\right)=\left(u_{1,1},\dots u_{1,m},u_{2,2},\dots,u_{2,m},\dots,u_{m,m}\right)$ and $\widetilde{\theta}=\left(\widetilde{\theta}^{1},\dots,\widetilde{\theta}^{m(m+1)/2}\right)=(\widetilde{u}_{1,1},\dots\widetilde{u}_{1,m},\widetilde{u}_{2,2},\dots,\widetilde{u}_{2,m},\dots,\widetilde{u}_{m,m})$, we shall compute the inner product defined by the Fisher metric, cf. (2.4). In the following, $\otimes$ denotes the Kronecker product, and, for an $m\times m$ matrix $A$, whose columns are $A_{1},A_{2},\dots,A_{m}$, we denote $\operatorname{vec}(A)\coloneqq\begin{pmatrix}A_{1}^{\mathsf{T}}&A_{2}^{\mathsf{T}}&\cdots&A_{m}^{\mathsf{T}}\end{pmatrix}^{\mathsf{T}}$ the $m^{2}$-dimensional vector formed by the concatenation of its columns. Moreover, if $A$ is symmetric, denote $\nu(A)\coloneqq(a_{1,1},\dots a_{1,m},a_{2,2},\dots,a_{2,m},\dots,a_{m,m})$. We denote $D_{m}$ the unique $m^{2}\times\frac{m(m+1)}{2}$ matrix that verifies $D_{m}\nu(A)=\operatorname{vec}(A)$, for any symmetric $A$ [43, § 7]. We thus have
$\langle U,V\rangle_{G(\xi)}$ $=\theta^{\mathsf{T}}G(\xi)\widetilde{\theta}$
@@
$=\frac{n}{2}\operatorname{tr}\left(\Sigma^{-1}U\Sigma^{-1}V\right)$
$=\frac{n}{2}\operatorname{vec}(U)(\Sigma^{-1}\otimes\Sigma^{-1})\operatorname{vec}(V)$
@@
where we have used that $\operatorname{tr}(ABCD)=\operatorname{vec}(D)^{\mathsf{T}}(A\otimes C^{\mathsf{T}})\operatorname{vec}(B^{\mathsf{T}})$, for $A$, $B$, $C$ and $D$ matrices such that the product $ABCD$ is defined and square [43, Lemma 3]. We can thus conclude that the Fisher matrix is
@@
$d_{\mathrm{FR}}\big{(}\Sigma_{1},\Sigma_{2}\big{)}=\sqrt{\frac{n}{2}}\left\|\log\left(\Sigma_{1}^{-1/2}\Sigma_{2}\Sigma_{1}^{-1/2}\right)\right\|_{F}=\sqrt{\frac{n}{2}\sum_{k=1}^{p}\left(\log\lambda_{k}\right)^{2}},$ (5.40)
where $\lambda_{k}$ are the eigenvalues of $\Sigma_{1}^{-1/2}\Sigma_{2}\Sigma_{1}^{-1/2}$, $\log$ denotes the matrix logarithm, and $\|A\|_{F}=\sqrt{\operatorname{tr}\left(AA^{\mathsf{T}}\right)}$ is the Frobenius norm. This also coincides with the standard metric of $P_{m}(\mathbb{R})$, when endowed with the matrix inner product $\langle A,B\rangle=\operatorname{tr}(A^{\mathsf{T}}B)$ [11, Chapter 6], up to the factor $\frac{n}{2}$, and is in fact related to the metric of the Siegel upper space [65] (see also [51, Appendix D]).
@@ 5.11 Inverse Wishart
where we have used that $\left|\frac{\mathrm{d}Y}{\mathrm{d}X}\right|=\det(X)^{-(m+1)}$ [43, § 12], which coincides with the p.d.f. of a Wishart distribution with parameter $\Phi$. Write $\sigma_{i,j}$ and $\phi_{i,j}$ the $(i,j)$-th entries of matrices $\Sigma$ and $\Phi$, respectively. Write $\xi=\left(\xi^{1},\dots,\xi^{m(m+1)/2}\right)=(\sigma_{1,1},\dots\sigma_{1,m},\sigma_{2,2},\dots,\sigma_{2,m},\dots,\sigma_{m,m})$, and $\theta=(\theta^{1},\dots,\theta^{m(m+1)/2})=(\phi_{1,1},\dots\phi_{1,m},\phi_{2,2},\dots,\phi_{2,m},\dots,\phi_{m,m})$. Denote $G_{W}(\theta)$ the Fisher matrix of a Wishart distribution (5.39) in coordinates $\theta$. In the following, $D_{m}$ is the matrix defined in the previous section, and $A^{+}$ denotes the Moore-Penrose inverse of a matrix $A$. Applying Proposition 2.3 yields
@@
$=\left(-D_{m}^{+}\left(\Phi^{-1}\otimes\Phi^{-1}\right)D_{m}\right)^{-\mathsf{T}}D_{m}^{\mathsf{T}}\left(\Phi^{-1}\otimes\Phi^{-1}\right)D_{m}\left(-D_{m}^{+}\left(\Phi^{-1}\otimes\Phi^{-1}\right)D_{m}\right)^{-1}$
$=\left(D_{m}^{+}\left(\Sigma\otimes\Sigma\right)D_{m}\right)^{-\mathsf{T}}D_{m}^{\mathsf{T}}\left(\Sigma\otimes\Sigma\right)D_{m}\left(D_{m}^{+}\left(\Sigma\otimes\Sigma\right)D_{m}\right)^{-1}$
$=\left(D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}\right)^{\mathsf{T}}D_{m}^{\mathsf{T}}\left(\Sigma\otimes\Sigma\right)D_{m}D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}$
$=D_{m}^{\mathsf{T}}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)\left(D_{m}^{+}\right)^{\mathsf{T}}D_{m}^{\mathsf{T}}\left(\Sigma\otimes\Sigma\right)D_{m}D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}$
$=D_{m}^{\mathsf{T}}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m},$
where we have used that $\frac{\mathrm{d}\Sigma}{\mathrm{d}\Phi}(\Phi)=-D_{m}^{+}\left(\Phi^{-1}\otimes\Phi^{-1}\right)D_{m}$ [43, § 12], $\left(D_{m}^{+}(\Sigma\otimes\Sigma)D_{m}\right)^{-1}=D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}$, and $D_{m}D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}=\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}$ [43, Lemma 11]. Therefore, the Fisher matrix is the same as for Wishart distributions in § (5.39), and the Fisher-Rao distance is
$d_{\mathrm{FR}}\big{(}\Sigma_{1},\Sigma_{2}\big{)}=\sqrt{\frac{n}{2}}\left\|\log\left(\Sigma_{1}^{-1/2}\Sigma_{2}\Sigma_{1}^{-1/2}\right)\right\|_{F}=\sqrt{\frac{n}{2}\sum_{k=1}^{p}\left(\log\lambda_{k}\right)^{2}},$ (5.41)
@@ 6 Product Distributions
where $G_{i}$ is the matrix form of the metric $g_{i}$, for $1\leq i\leq m$. Let $d_{i}$ denote the geodesic distance in $M_{i}$. Then, the geodesic distance $d$ in $(M,g)$ is given by a Pythagorean formula [24, Prop. 1], [53]:
@@
$=\sqrt{\sum_{i=1}^{n}(\beta_{i}+1)\left[\operatorname{arctanh}\left(\sqrt{\frac{\beta_{i}(\mu_{i}-\mu^{\prime}_{i})^{2}\Gamma(2-1/\beta_{i})+(\beta_{i}+1)(\sigma_{i}-\sigma^{\prime}_{i})^{2}\Gamma(1+1/\beta_{i})}{\beta_{i}(\mu_{i}-\mu^{\prime}_{i})^{2}\Gamma(2-1/\beta_{i})+(\beta_{i}+1)(\sigma^{\prime}_{i}+\sigma^{\prime}_{i})^{2}\Gamma(1+1/\beta_{i})}}\right)\right]^{2}}.$ (6.4)

### v3

@@ 4.9 Power function
    which is negative, but not constant, and where $\psi^{(m)}(x)\coloneqq\frac{\mathrm{d}^{m+1}}{\mathrm{d}x^{m+1}}\log\Gamma(x)$ denotes the polygamma function of order $m$. Bounds for the Fisher–Rao distance in this manifold have been studied in [64].
@@
    $\kappa$ $=\frac{\psi^{(2)}(\alpha)\psi^{(2)}(\beta)\psi^{(2)}(\alpha+\beta)}{4\left(\psi^{(1)}(\alpha)\psi^{(1)}(\beta)-\psi^{(1)}(\alpha+\beta)[\psi^{(1)}(\alpha)+\psi^{(1)}(\beta)]\right)^{2}}$
    ×$\quad\times\left(\frac{\psi^{(1)}(\alpha)}{\psi^{(2)}(\alpha)}+\frac{\psi^{(1)}(\beta)}{\psi^{(2)}(\beta)}-\frac{\psi^{(1)}(\alpha+\beta)}{\psi^{(2)}(\alpha+\beta)}\right)$
@@
    which is negative, but not constant too. In fact, more generally, the sectional curvature of the statistical manifold of Dirichlet distributions (which are the multivariate generalisation of Beta distributions) is negative [65, Thm. 6].
@@ 4.10 Wishart
defined for $X\in P_{m}(\mathbb{R})$, and characterised by $\Sigma\in P_{m}(\mathbb{R})$, where $\Gamma_{m}(z)\coloneqq\pi^{\frac{m(m-1)}{4}}\prod_{j=1}^{m}\Gamma\left(z+\frac{1-j}{2}\right)$ denotes the multivariate Gamma function. Note that, for fixed $m,n$, the associated $\left(\frac{m(m+1)}{2}\right)$-dimensional statistical manifold $\mathcal{S}$ is in correspondence with the cone $P_{m}(\mathbb{R})$ of symmetric positive-definite matrices via the bijection $\iota\colon P_{m}(\mathbb{R})\to\mathcal{S}$, given by $\Sigma\mapsto\frac{{(\det X)}^{(n-m-1)/2}\exp\left(-\frac{1}{2}\operatorname{tr}(\Sigma^{-1}X)\right)}{2^{nm/2}(\det\Sigma)^{n/2}\,\Gamma_{m}(n/2)}$. Denoting $\sigma_{i,j}$ the $(i,j)$-th entry of the matrix $\Sigma$, we can write the parameter vector as $\xi=\left(\xi^{1},\dots,\xi^{m(m+1)/2}\right)=(\sigma_{1,1},\dots\sigma_{1,m},\sigma_{2,2},\dots,\sigma_{2,m},\dots,\sigma_{m,m})$. We then have $\partial_{i}\ell\coloneqq\partial_{{i}}\ell(\Sigma)=\frac{1}{2}\operatorname{tr}\left(\Sigma^{-1}X\Sigma^{-1}{(\partial_{i}\Sigma)}\right)-\frac{n}{2}\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma)}\right)$ and $\partial_{j}\partial_{i}\ell=-\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma)}\Sigma^{-1}{(\partial_{j}\Sigma)}\Sigma^{-1}X\right)+\frac{n}{2}\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma)}\Sigma^{-1}{(\partial_{j}\Sigma)}\right)$, where the derivative in $\partial_{i}\Sigma$ is taken entry-wise. The elements of the Fisher metric are then
$g_{ij}(\xi)$ $=-\mathbb{E}\left[\partial_{j}\partial_{i}\ell\right]$
$=-\mathbb{E}\left[-\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma)}\Sigma^{-1}{(\partial_{j}\Sigma)}\Sigma^{-1}X\right)+\frac{n}{2}\operatorname{tr}\left(\Sigma^{-1}{(\partial_{i}\Sigma})\Sigma^{-1}{(\partial_{j}\Sigma)}\right)\right]$
@@
In view of the bijection $\iota$, the tangent space $T_{p_{\xi}}\mathcal{S}$ can be identified with $H_{m}$, the set of $m\times m$ symmetric matrices [67, Chapter 6]. Given two matrices $U,\ \widetilde{U}\in H_{m}$, parametrized as $\theta=\left(\theta^{1},\dots,\theta^{m(m+1)/2}\right)=\left(u_{1,1},\dots u_{1,m},u_{2,2},\dots,u_{2,m},\dots,u_{m,m}\right)$ and $\widetilde{\theta}=\left(\widetilde{\theta}^{1},\dots,\widetilde{\theta}^{m(m+1)/2}\right)=(\widetilde{u}_{1,1},\dots\widetilde{u}_{1,m},\widetilde{u}_{2,2},\dots,\widetilde{u}_{2,m},\dots,\widetilde{u}_{m,m})$, we shall compute the inner product defined by the Fisher metric, cf. (1.4). In the following, $\otimes$ denotes the Kronecker product, and, for an $m\times m$ matrix $A$, whose columns are $A_{1},A_{2},\dots,A_{m}$, we denote $\operatorname{vec}(A)\coloneqq\begin{pmatrix}A_{1}^{\mathsf{T}}&A_{2}^{\mathsf{T}}&\cdots&A_{m}^{\mathsf{T}}\end{pmatrix}^{\mathsf{T}}$ the $m^{2}$-dimensional vector formed by the concatenation of its columns. Moreover, if $A$ is symmetric, denote $\nu(A)\coloneqq(a_{1,1},\dots a_{1,m},a_{2,2},\dots,a_{2,m},\dots,a_{m,m})$. We denote $D_{m}$ the unique $m^{2}\times\frac{m(m+1)}{2}$ matrix that verifies $D_{m}\nu(A)=\operatorname{vec}(A)$, for any symmetric $A$ [68, § 7]. We thus have
$\langle U,\widetilde{U}\rangle_{G(\xi)}$ $=\theta^{\mathsf{T}}G(\xi)\widetilde{\theta}$
@@
$=\frac{n}{2}\operatorname{tr}\left(\Sigma^{-1}U\Sigma^{-1}\widetilde{U}\right)$
$=\frac{n}{2}\operatorname{vec}(U)(\Sigma^{-1}\otimes\Sigma^{-1})\operatorname{vec}(\widetilde{U})$
@@
where we have used that $\operatorname{tr}(ABCD)=\operatorname{vec}(D)^{\mathsf{T}}(A\otimes C^{\mathsf{T}})\operatorname{vec}(B^{\mathsf{T}})$, for $A$, $B$, $C$ and $D$ matrices such that the product $ABCD$ is defined and square [68, Lemma 3]. We can thus conclude that the Fisher matrix is
@@
$d_{\mathrm{FR}}\big{(}\Sigma_{1},\Sigma_{2}\big{)}=\sqrt{\frac{n}{2}}\left\|\log\left(\Sigma_{1}^{-1/2}\Sigma_{2}\Sigma_{1}^{-1/2}\right)\right\|_{F}=\sqrt{\frac{n}{2}\sum_{k=1}^{m}\left(\log\lambda_{k}\right)^{2}},$ (4.40)
where $\lambda_{k}$ are the eigenvalues of $\Sigma_{1}^{-1/2}\Sigma_{2}\Sigma_{1}^{-1/2}$, $\log$ denotes the matrix logarithm, and $\|A\|_{F}=\sqrt{\operatorname{tr}\left(AA^{\mathsf{T}}\right)}$ is the Frobenius norm. This metric also coincides with the standard metric of $P_{m}(\mathbb{R})$, when endowed with the matrix inner product $\langle A,B\rangle=\operatorname{tr}(A^{\mathsf{T}}B)$ [67, Chapter 6], up to the factor $\frac{n}{2}$, and is in fact related to the metric of the Siegel upper space [69] (see also [37, Appendix D]). Note that when $\Sigma$ is restricted to be diagonal this distance is, up to a factor $\sqrt{n}$, the product distance between univariate Gaussian distributions with fixed mean—see Example 5.2 in Section 5.
@@ 4.11 Inverse Wishart
where we have used that $\left|\frac{\mathrm{d}Y}{\mathrm{d}X}\right|=\det(X)^{-(m+1)}$ [68, § 12], which coincides with the p.d.f. of a Wishart distribution with parameter $\Phi$. Write $\sigma_{i,j}$ and $\phi_{i,j}$ the $(i,j)$-th entries of matrices $\Sigma$ and $\Phi$, respectively. Write $\xi=\left(\xi^{1},\dots,\xi^{m(m+1)/2}\right)=(\sigma_{1,1},\dots\sigma_{1,m},\sigma_{2,2},\dots,\sigma_{2,m},\dots,\sigma_{m,m})$, and $\theta=(\theta^{1},\dots,\theta^{m(m+1)/2})=(\phi_{1,1},\dots\phi_{1,m},\phi_{2,2},\dots,\phi_{2,m},\dots,\phi_{m,m})$. Denote $G_{W}(\theta)$ the Fisher matrix of a Wishart distribution (4.39) in coordinates $\theta$. In the following, $D_{m}$ is the matrix defined in the previous section, and $A^{+}$ denotes the Moore-Penrose inverse of a matrix $A$. Applying Proposition 1.2 yields
@@
$=\frac{n}{2}\left(-D_{m}^{+}\left(\Phi^{-1}\otimes\Phi^{-1}\right)D_{m}\right)^{-\mathsf{T}}D_{m}^{\mathsf{T}}\left(\Phi^{-1}\otimes\Phi^{-1}\right)D_{m}\left(-D_{m}^{+}\left(\Phi^{-1}\otimes\Phi^{-1}\right)D_{m}\right)^{-1}$
$=\frac{n}{2}\left(D_{m}^{+}\left(\Sigma\otimes\Sigma\right)D_{m}\right)^{-\mathsf{T}}D_{m}^{\mathsf{T}}\left(\Sigma\otimes\Sigma\right)D_{m}\left(D_{m}^{+}\left(\Sigma\otimes\Sigma\right)D_{m}\right)^{-1}$
$=\frac{n}{2}\left(D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}\right)^{\mathsf{T}}D_{m}^{\mathsf{T}}\left(\Sigma\otimes\Sigma\right)D_{m}D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}$
$=\frac{n}{2}D_{m}^{\mathsf{T}}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)\left(D_{m}^{+}\right)^{\mathsf{T}}D_{m}^{\mathsf{T}}\left(\Sigma\otimes\Sigma\right)D_{m}D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}$
$=\frac{n}{2}D_{m}^{\mathsf{T}}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m},$
where we have used that $\frac{\mathrm{d}\Sigma}{\mathrm{d}\Phi}(\Phi)=-D_{m}^{+}\left(\Phi^{-1}\otimes\Phi^{-1}\right)D_{m}$ [68, § 12], $\left(D_{m}^{+}(\Sigma\otimes\Sigma)D_{m}\right)^{-1}=D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}$, and $D_{m}D_{m}^{+}\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}=\left(\Sigma^{-1}\otimes\Sigma^{-1}\right)D_{m}$ [68, Lemma 11]. Therefore, the Fisher matrix is the same as for Wishart distributions in § (4.39), and the Fisher-Rao distance is
$d_{\mathrm{FR}}\big{(}\Sigma_{1},\Sigma_{2}\big{)}=\sqrt{\frac{n}{2}}\left\|\log\left(\Sigma_{1}^{-1/2}\Sigma_{2}\Sigma_{1}^{-1/2}\right)\right\|_{F}=\sqrt{\frac{n}{2}\sum_{k=1}^{m}\left(\log\lambda_{k}\right)^{2}},$ (4.41)
@@ 5 Product Distributions
where $G_{i}$ is the matrix form of the metric $g_{i}$, for $1\leq i\leq m$. Let $d_{i}$ denote the geodesic distance in $M_{i}$. Then, the geodesic distance $d$ in $(M,g)$ is given by a Pythagorean formula [71, Prop. 1], [72]:
@@
$=\sqrt{\sum_{i=1}^{n}(\beta_{i}+1)\left[\operatorname{arctanh}\left(\sqrt{\frac{\beta_{i}(\mu_{i}-\mu^{\prime}_{i})^{2}\Gamma(2-\frac{1}{\beta_{i}})+(\beta_{i}+1)(\sigma_{i}-\sigma^{\prime}_{i})^{2}\Gamma(1+\frac{1}{\beta_{i}})}{\beta_{i}(\mu_{i}-\mu^{\prime}_{i})^{2}\Gamma(2-\frac{1}{\beta_{i}})+(\beta_{i}+1)(\sigma^{\prime}_{i}+\sigma^{\prime}_{i})^{2}\Gamma(1+\frac{1}{\beta_{i}})}}\right)\right]^{2}}.$ (5.4)